OpenStudy (jannine24):

need help solving and graphing following functions: 1. y=(-3/4)x + 5 2.y= x^2-4 3. y= -3x^2 -6x -5

OpenStudy (jannine24):

i know the first one is a line and other 2 are parboles

OpenStudy (anonymous):

yes, you can use a table coordinate system for number 1

OpenStudy (anonymous):

just substitute x = 0,1,2,3,4, and you get your line, easy.

OpenStudy (jannine24):

for 1 i have y interept of 5

OpenStudy (anonymous):

second one, it's a perfect square x^2 - 4 = (x-2)(x+2)

OpenStudy (jannine24):

and xintercept of 6 2/3

OpenStudy (jannine24):

the second one is a parbole correct

OpenStudy (jannine24):

so i need ot find the vertex first

OpenStudy (anonymous):

Yep. so you can graph it now, since (x-2)(x+2)= y. just plot in points x = 0,1,2,3,4

OpenStudy (anonymous):

and once you see the graph, you'll get the vertex. Visually seeing it first can make it easier.

OpenStudy (jannine24):

i rather use the formula

OpenStudy (anonymous):

do you have the formula?

OpenStudy (jannine24):

i think this is it -b/2(a) right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-b/2a, then use this number and plug it in the original f(x) to get the y coordinate

OpenStudy (jannine24):

amoondarya ty can u show me how u solved it

OpenStudy (jannine24):

they r three seperate problems

OpenStudy (jannine24):

dpalnc any ideas

OpenStudy (jannine24):

y= (-3/4)x +5 to have a y intercept put 0 in for x and get 5 so y intercept is (0,5)

OpenStudy (jannine24):

the xintercept would be 0 for y and you get 6 2/3 so it would be (6 2/3 , 0) for x intercept

OpenStudy (jannine24):

then a thrid point could be 4 for x then get y= (-3/4)4 + 5 so it would be y=-3 + 5 so y=2 so thrid point is (4,2)

OpenStudy (amoodarya):